# Corresponding author(s)

### v4

@@ Abstract
DeepSeek-R1, known for its low training cost and exceptional reasoning capabilities, has achieved state-of-the-art performance on various benchmarks. However, detailed evaluations for DeepSeek Series models from the perspective of real-world applications are lacking, making it challenging for users to select the most suitable DeepSeek models for their specific needs. To address this gap, we conduct a systematic evaluation of the DeepSeek-V3, DeepSeek-R1, DeepSeek-R1-Distill-Qwen series, DeepSeek-R1-Distill-Llama series, their corresponding 4-bit quantized models, and the reasoning model QwQ-32B using the enhanced A-Eval benchmark, A-Eval-2.0. Through a comparative analysis of original instruction-tuned models and their distilled counterparts, we investigate how reasoning enhancements impact performance across diverse practical tasks. Our results show that reasoning-enhanced models, while generally powerful, do not universally outperform across all tasks, with performance gains varying significantly across tasks and models. To assist users in model selection, we quantify the capability boundary of DeepSeek models through performance tier classifications. Based on the quantification results, we develop a model selection handbook that clearly illustrates the relation among models, their capabilities and practical applications. This handbook enables users to select the most cost-effective models without efforts, ensuring optimal performance and resource efficiency in real-world applications. It should be noted that, despite our efforts to establish a comprehensive, objective, and authoritative evaluation benchmark, the selection of test samples, characteristics of data distribution, and the setting of evaluation criteria may inevitably introduce certain biases into the evaluation results. We will continuously optimize the evaluation benchmarks and periodically update this paper to provide more comprehensive and accurate evaluation results. Please refer to the latest version of the paper for the most current results and conclusions.
Keywords:
DeepSeek application-driven evaluation reasoning enhancements capability boundary model selection handbook
@@ 1 Introduction
To address the gap, we first create A-Eval-2.0 by modifying some QAs in A-Eval [10], an application-driven evaluation benchmark comprising 678 human-curated question-answer (QA) pairs across five major task categories (Text Understanding, Information Extraction, Text Generation, Logical Reasoning, and Task Planning) and 27 subcategories. Then, we evaluate DeepSeek models on A-Eval-2.0. Through this evaluation, we extend existing evaluations and provide actionable insights into how reasoning enhancements influence model performance across various practical tasks. These results and findings help guide users in selecting the most suitable DeepSeek models for their specific needs, enabling cost-effective deployment in real-world applications.
@@ 2.2 Dataset
The A-Eval [10] benchmark, designed for practical application scenarios, quantifies the capability boundaries of LLMs through 678 manually curated QA pairs across three difficulty level, five major categories and 27 subcategories. To accommodate the enhanced reasoning capabilities of the DeepSeek series models, we refine the sub-tasks and QAs of Task Planning category in A-Eval, thereby constructing the evaluation dataset A-Eval-2.0 utilized in this evaluation. We use A-Eval-2.0’s dataset as the evaluation dataset to analyze the performance of DeepSeek series models in practical applications. On one hand, the evaluation on A-Eval-2.0 enables researchers to assess how the reasoning enhancements in DeepSeek-R1 improve model performance across various practical scenarios. On the other hand, the evaluation results assist users in selecting the most cost-effective models for their specific application needs.
@@ 2.3 Evaluation Process
We follow the automatic and manual scoring methods of A-Eval, and the automatic scoring process includes:
@@
Scoring: Combine the prompt and triplet$\left({{Q_{i}},{A_{i}},{P_{i}}}\right)$, input them into the scoring model, and obtain the scoring output ${{S_{i}}}$ between 0 and 100. We use A-Eval’s scoring prompt and a greater LLM, Qwen2.5-72B-Instruct as the scoring model.
@@ 3 Results and Discussion
In this section, we first present the overall average performance of 22 models across all data in A-Eval-2.0 (Section 3.1). Subsequently, we analyze the models’ performance from three perspectives: task category (Section 3.2), sub-task category (Section 3.3), and model group (Section 3.4). Following the A-Eval methodology, we visualize the comparative evaluation results using line charts (Section 3.5). Then, we statistic performance of models on datasets with three difficulty levels—easy, medium, and hard—and discuss models’ performance to variations in question difficulty (Section 3.6). Finally, we quantify models’ capability boundary through a tier classification table and develop a model selection handbook to assist users in selecting the most cost-effective models for their specific requirements (Section 3.7).
@@ 3.5 Line Chart Results
Figure 4: Line chart for model selection on five major tasks.
Figure 4 illustrates the scores of the 22 evaluated models on each major task, as well as their average scores across all tasks. This comprehensive overview allows users to quickly identify models that excel in specific major task types or demonstrate strong overall performance.
In Figure 5, we further break down the performance of the 22 models by showing their average scores within each of the five major task categories and their corresponding subcategories. These detailed visualizations enable users to assess model capabilities at a granular level, ensuring a precise match between model strengths and task requirements.
Figure 5: Line chart for model selection on five major tasks and corresponding subtasks.
@@ 3.6 Performance Comparison By Difficulty Level
3.7 Model Selection Guidance for Users
To assist users in selecting the most suitable model based on their specific application requirements, we classify model performance into five tiers (A+, A, B, C, D) based on their scores in the five major task categories, resulting in the capability boundary quantification table (Table 2).
@@ 3.7 Model Selection Guidance for Users
To further simplify the model selection process for users without extensive knowledge of LLMs, we comb comprehensive real-world applications and develop the model selection handbook that is shown in Figure 8.
Figure 8: The model selection handbook.
The model selection handbook illustrates the relation among models, their capabilities, and practical applications, enabling users to select the most cost-effective model without any effort.
It should be noted that any evaluation benchmark inherently has certain limitations, including the selection of test samples, the characteristics of data distribution, and the setting of evaluation criteria. Therefore, our evaluation results may not be the most objective, comprehensive, or accurate. As we continue to improve this evaluation work, we recommend referring to the latest version of the paper for the most recent experimental results and conclusions.

### v5

@@ Abstract
DeepSeek-R1, known for its low training cost and exceptional reasoning capabilities, has achieved state-of-the-art performance on various benchmarks. However, detailed evaluations for DeepSeek Series models from the perspective of real-world applications are lacking, making it challenging for users to select the most suitable DeepSeek models for their specific needs. To address this gap, we presents the first comprehensive evaluation of the DeepSeek and its related models (including DeepSeek-V3, DeepSeek-R1, DeepSeek-R1-Distill-Qwen series, DeepSeek-R1-Distill-Llama series, their corresponding 4-bit quantized models, and the reasoning model QwQ-32B) using our enhanced A-Eval benchmark, A-Eval-2.0. Our systematic analysis reveals several key insights: (1) Given identical model architectures and training data, larger parameter models demonstrate superior performance, aligning with the scaling law. However, smaller models may achieve enhanced capabilities when employing optimized training strategies and higher-quality data; (2) Reasoning-enhanced model show significant performance gains in logical reasoning tasks but may underperform in text understanding and generation tasks; (3) As the data difficulty increases, distillation or reasoning enhancements yield higher performance gains for the models. Interestingly, reasoning enhancements can even have a negative impact on simpler problems; (4) Quantization impacts different capabilities unevenly, with significant drop on logical reasoning and minimal impact on text generation. Based on these results and findings, we design an model selection handbook enabling users to select the most cost-effective models without efforts. It should be noted that, despite our efforts to establish a comprehensive, objective, and authoritative evaluation benchmark, the selection of test samples, characteristics of data distribution, and the setting of evaluation criteria may inevitably introduce certain biases into the evaluation results. We will continuously optimize the evaluation benchmarks and periodically update this paper to provide more comprehensive and accurate evaluation results. Please refer to the latest version of the paper for the most current results and conclusions.
@@ 1 Introduction
To address the gap, we first create A-Eval-2.0 by modifying some QAs in our previous A-Eval [10], an application-driven evaluation benchmark comprising 678 human-curated question-answer (QA) pairs across five major task categories (Text Understanding, Information Extraction, Text Generation, Logical Reasoning, and Task Planning) and 27 subcategories. Then, we evaluate DeepSeek models on A-Eval-2.0. Through this evaluation, we extend existing evaluations and provide actionable insights into how reasoning enhancements influence model performance across various practical tasks. These results and findings help guide users in selecting the most suitable DeepSeek models for their specific needs, enabling cost-effective deployment in real-world applications.
Based on these results and findings, we contribute: (a) The open-sourced A-Eval-2.0 benchmark, including dataset and an evaluation toolkit; (b) The capability boundary quantification results for DeepSeek models; (c) An model selection handbook enabling users to select the most cost-effective models without efforts. Our work provides both methodological contributions to LLM evaluation and practical guidance for real-world applications.
@@ 2.2 Dataset
The A-Eval benchmark [10] was originally designed to assess the capability boundaries of large language models (LLMs) in practical application scenarios. It consists of 678 manually curated question-answer pairs spanning three difficulty levels, five major categories, and 27 subcategories. While comprehensive for its initial purpose, A-Eval was primarily designed for evaluating conventional instruction-tuned models across different parameter scales.
However, the original benchmark has several limitations in the context of modern LLM development:
1. It does not account for specialized model variants like distilled, quantized, or reasoning-optimized architectures.
2. Its evaluation granularity is insufficient for complex reasoning and planning tasks.
3. The difficulty progression of data lacks precision in assessing advanced cognitive capabilities.
To address these limitations and better evaluate the enhanced reasoning capacities of DeepSeek models, we introduce A-Eval-2.0. This enhanced version maintains the core structure of A-Eval while incorporating critical improvements in two key aspects:
Extended Task Planning Subcategories: We add "Scheme Optimization" as a new subcategory under Task Planning category and make 40 new test QA pairs, focusing on evaluating models’ ability to execute multi-step plans under constraints and optimize the plans according to dynamic feedback. This sub-task can more fully test the complex planning ability of the reasoning models.
Refined Logical Reasoning and Task Planning Categories: We refine Logical Reasoning and Task Planning ability into Primary, Intermediate and Advanced levels and make clear definitions:
Primary Logical Reasoning: Processes well-defined logical verification or contradiction detection tasks, typically accomplished through simple rules and single-step causal inference.
Intermediate Logical Reasoning: Accomplishes multi-factor, multi-step causal reasoning within specific domains by processing conventional constraints and establishing localized logical chains.
Advanced Logical Reasoning: Addresses domain-specific complex reasoning challenges requiring professional knowledge integration for comprehensive systemic deduction, backtracking, and refinement, ultimately generating professional-grade decisions or solutions.
Primary Task Planning: Formulates linearly decomposable deterministic task flows through standardized breakdown of finite steps, executing basic commonsense planning with either explicit parallel branches or no parallelism, without requiring dynamic adjustment.
Intermediate Task Planning: Designs multi-constrained comprehensive solutions for single entities by balancing resources, priorities, and personalized needs, while handling common limited variants and optimizing execution paths.
Advanced Task Planning: Constructs general-purpose intelligent agents capable of professional-grade, system-level planning in dynamic multi-agent environments, integrating real-time environmental variables to adapt execution strategies.
We use A-Eval-2.0’s dataset as the evaluation dataset to analyze the performance of DeepSeek series models in practical applications. On one hand, the evaluation on A-Eval-2.0 enables researchers to assess how the reasoning enhancements in DeepSeek-R1 improve model performance across various practical scenarios. On the other hand, the evaluation results assist users in selecting the most cost-effective models for their specific application needs.
@@ 2.3 Evaluation Process
We follow the automatic scoring pipeline of original A-Eval, and the automatic scoring process includes:
@@
Scoring: Combine the prompt and triplet$\left({{Q_{i}},{A_{i}},{P_{i}}}\right)$, input them into the scoring model, and obtain the scoring output ${{S_{i}}}$ between 0 and 100.
To better evaluate the enhanced reasoning capacities of DeepSeek models, we optimize the scoring details in several aspects:
Advanced Scoring Model: We employ DeepSeek-R1 as the scoring model, which demonstrates superior judgment capabilities compared to original instruction-tuned model.
Optimized Scoring Prompt: To better assess the capabilities of instruction-tuned models, distilled models, reasoning models and quantized models, we refine the scoring prompt.
Two-Phase Expert Review: To ensure the reliability of the evaluation results, our evaluation expert team conducts a two-round review of the automatic scoring results:
Phase 1. Independent Reviewing: Firstly, $N$ evaluation experts independently assess each automated scoring tuple $\left({{Q_{i}},{A_{i}},{P_{i}},{S_{i}}}\right)$ after aligning expert scoring standards. Each expert may output the manual score $E_{i}^{k}={S_{i}}$ if the automated score ${{S_{i}}}$ is deemed reasonable or provide the manual score $E_{i}^{k}\neq{S_{i}}$, where $k=\left\{{1,2,....,N}\right\}$. For each tuple, we then calculate the consensus ratio: $C{R_{i}}={1\over N}\sum\limits_{k=1}^{N}{{\rm I}\left({E_{i}^{k}={S_{i}}}\right)}$ where ${\rm I}$ denotes the indicator function. Finally, ${S_{i}}$ is considered reliable if $C{R_{i}}\geq\tau$ (we set $N=5$ and $\tau=0.8$), indicating at least 80% expert agreement with the automated score ${S_{i}}$.
Phase 2. Consensus Reconciliation: For tuples failing to meet the consensus threshold ($C{R_{i}}<\tau$), the evaluation team conducts a focused group discussion. Following discussion, each expert submits a revised score $E_{i}^{k*}$. The final revised score $S_{i}^{*}$ is computed using averaging method that discards the highest and lowest scores.
This two-phase review process ensures both independent assessment and collective wisdom are incorporated into the final evaluation results. The optimized scoring prompt is as follows:
Scoring Prompt Template
@@ 3 Results and Discussion
In this section, we first present the overall average performance of 22 models across all data in A-Eval-2.0 (Section 3.1). Subsequently, we analyze the models’ performance from three perspectives: task category (Section 3.2), sub-task category (Section 3.3), and model group (Section 3.4). Following the A-Eval methodology, we visualize the comparative evaluation results using line charts (Section 3.5). Then, we statistic performance of models on datasets with three difficulty levels—easy, medium, and hard—and discuss models’ performance to variations in question difficulty (Section 3.6). Then, we quantify models’ capability boundary through a tier classification table, illustrating performance levels of models on different capabilities (Section 3.7). Finally, develop a model selection handbook to assist users in selecting the most cost-effective models for their specific requirements without efforts (Section 3.8).
@@ 3.5 Line Chart Results
Figure 4: Line charts for evaluation performance. (a) Line chart on five major tasks. (b) - (f) Line charts on five major tasks and corresponding subtasks
Figure 4 (a) illustrates the scores of the 22 evaluated models on each major task, as well as their average scores across all tasks. This comprehensive overview allows users to quickly identify models that excel in specific major task types or demonstrate strong overall performance.
In Figure 4 (b) to Figure 4 (f), we further break down the performance of the 22 models by showing their average scores within each of the five major task categories and their corresponding subcategories. These detailed visualizations enable users to assess model capabilities at a granular level, ensuring a precise match between model strengths and task requirements.
@@ 3.6 Performance Comparison By Difficulty Level
3.7 Model Performance Tier Classifications
We classify model performance into five tiers (A+, A, B, C, D) based on their scores in the five major task categories, resulting in the capability boundary quantification table (Table 2).
@@ 3.7 Model Performance Tier Classifications
3.8 Model Selection Notebook for Users
To assist users in selecting the most suitable model based on their specific application requirements, we develop a model selection notebook consisting of four components.
1. LLM Capability Definitions and Application Scenarios. As shown in Figure 7, we systematically categorize a substantial number of real-world application scenarios and analyze the specific model capabilities required for each.
Figure 7: LLM capability definitions and application scenarios, including 15 well-defined model capabilities along with their corresponding 105 application scenarios.
2. LLM “Parameter Scale-Capability-Application Scenarios” Relation. As shown in Figure 8, we analyze the relationship among parameter scale, capability, and application based on the evaluation results and the model performance tier classifications in Table 2.
Figure 8: LLM “Parameter Scale-Capability-Application” relation. The figure illustrates the correlations among models, their capabilities, and practical applications. For each capability, the most cost-effective model is indicated by green lines, while other models with higher computational costs are shown in red.
3. Supported Application Scenarios for Each Model. As shown in Figure 9, we clearly demonstrate the specific application scenarios that each model can successfully accomplish.
Figure 9: Supported application scenarios of each model.
4. Cost-Effective Model Recommendations for Target Application Scenarios. As illustrated in Figure 10, to simplify the model selection process for users without extensive knowledge of LLMs, we provide cost-effective model recommendations for each application scenario, along with detailed hardware deployment requirements.
Figure 10: Cost-Effective model recommendations for target application scenarios.
Using our model selection notebook, users can efficiently identify the most cost-effective models, ensuring optimal performance and resource efficiency in real-world applications.
It should be noted that, despite our efforts to establish a comprehensive, objective, and authoritative evaluation benchmark, the selection of test samples, characteristics of data distribution, and the setting of evaluation criteria may inevitably introduce certain biases into the evaluation results. We will continuously optimize the evaluation benchmarks and periodically update this paper to provide more comprehensive and accurate evaluation results. Please refer to the latest version of the paper for the most current results and conclusions.